# Back on Track: Bundle Adjustment for Dynamic Scene Reconstruction

### v1

@@ Abstract
Traditional SLAM systems, which rely on bundle adjustment, struggle with highly dynamic scenes commonly found in casual videos. Such videos entangle the motion of dynamic elements, undermining the assumption of static environments required by traditional systems. Existing techniques either filter out dynamic elements or model their motion independently. However, the former often results in incomplete reconstructions, whereas the latter can lead to inconsistent motion estimates. Taking a novel approach, this work leverages a 3D point tracker to separate the camera-induced motion from the observed motion of dynamic objects. By considering only the camera-induced component, bundle adjustment can operate reliably on all scene elements as a result. We further ensure depth consistency across video frames with lightweight post-processing based on scale maps. Our framework combines the core of traditional SLAM – bundle adjustment – with a robust learning-based 3D tracker front-end. Integrating motion decomposition, bundle adjustment and depth refinement, our unified framework, BA-Track, accurately tracks the camera motion and produces temporally coherent and scale-consistent dense reconstructions, accommodating both static and dynamic elements. Our experiments on challenging datasets reveal significant improvements in camera pose estimation and 3D reconstruction accuracy. Project website: https://wrchen530.github.io/projects/batrack/
@@ 1 Introduction
Figure 2: Dynamic scene reconstruction results on Shibuya [30], DAVIS [19], and Aria Everyday Activities [26] datasets. By leveraging sparse dynamic SLAM and global refinement, BA-Track achieves consistent dense 3D reconstructions across diverse dynamic scenes.
@@
With the increasing prevalence of casual videos, reconstructing scenes containing moving objects has become essential for applications like augmented reality and robotics. Traditional methods for structure-from-motion (SfM) [1] and simultaneous localization and mapping (SLAM) [12] excel in recovering camera poses and scene geometry. However, they leverage the epipolar constraint and thus apply only to static environments. Dynamic scenes with moving objects introduce ambiguities that violate the epipolar constraint. Resolving these ambiguities by analytical constraints or statistical outlier filtering alone is infeasible due to the complexity and diversity of real-world object motion. This challenge motivates a hybrid approach that combines traditional optimization with deep priors. Existing hybrid methods often address dynamic elements by detecting and removing them from optimization, leading to incomplete reconstructions or missing details associated with moving objects [2, 48]. Another approach is to rely on geometric cues like monocular depth model for per-frame reconstruction, but requires additional optimization to achieve a globally aligned 3D reconstruction [21]. This proves challenging in practice due to the inconsistency between estimated camera pose and depth priors across different frames, which can result in misaligned depth maps when integrating them into a coherent 3D scene.
@@
We propose BA-Track, a framework that jointly reconstructs static and dynamic scene geometry from casual video sequences (see Fig. 1). Following traditional SfM and SLAM approaches, BA-Track extends the time-honored bundle adjustment (BA) [39] to dynamic scene reconstruction. The novel insight is to integrate a learning-based 3D tracker to decouple the motion of dynamic elements from the observed motion. By singling out the camera-induced motion component of dynamic objects, BA-Track reconstructs 3D trajectories of dynamic points as if they were stationary relative to their local reference frame. As a result, the epipolar constraint becomes applicable to all points, whether dynamic or static, enabling robust operation of bundle adjustment.
@@
To effectively decouple motion, we incorporate monocular depth priors with a 3D tracker. The tracker learns valuable 3D priors that help distinguishing between camera- and object-based motion. While such learning-based priors prove extremely helpful for reconstructing dynamic scenes, the scale of the depth maps exhibits temporal and spatial inconsistency. To address this issue, we develop a global refinement module, which leverages the accurate but sparse geometry from BA to refine the dense depth maps. As illustrated in Fig. 2, BA-Track achieves compelling reconstruction quality of both static and dynamic scene elements.
@@
In summary, our framework consists of three components: (1) A robust 3D tracking front-end which decouples the camera-induced motion from the observed (total) motion; (2) a bundle adjustment back-end for accurate pose and depth estimation, which leverages the camera-induced motion for all points; and (3) global refinement, to achieve dense, scale-consistent depth. Our learnable pipeline extends epipolar geometry to dynamic scene reconstruction, leveraging the power of established optimization techniques. Our experiments demonstrate significant improvements in camera motion estimation and 3D scene reconstruction, even in challenging scenes.
@@ 2 Related Work
Point Tracking.   Point tracking techniques estimate 2D positions of one or more query points across a sequence of images. Revisiting the idea of video representation as moving particles [31], PIPs [15] takes a learning-based approach with a feed-forward network, which iteratively refines point trajectories using multi-scale local features. TAP-Net [10] takes a different direction by leveraging global correlations to directly compute trajectories, thereby eliminating the need for iterative refinement. Subsequent research has explored various techniques to improve accuracy and efficiency, including inter-track relationships [17], temporal refinements [11], nearest-neighbor interpolation [22] and bidirectional 4D correlation [8]. Moving beyond 2D tracking, recent methods, such as SpatialTracker [44] and SceneTracker [2], elevate the task into 3D space by incorporating depth priors. Our approach extends 3D tracking by introducing a novel motion decoupling strategy that separates camera motion from object motion, enabling more effective integration with bundle adjustment.
@@
Monocular Visual Odometry.   Monocular visual odometry (VO) estimates camera motion using RGB image sequences. Traditional methods can be divided into feature-based and direct approaches. Feature-based methods detect and match keypoints across frames and use epipolar geometry with bundle adjustment for pose estimation [9, 27]. In contrast, direct methods bypass keypoints and optimize photometric consistency between frames, leveraging information from the entire image [12, 14, 13]. Recent advances have shifted towards learning-based approaches that employ neural networks to enhance feature extraction and pose estimation [5, 36, 37, 35, 43, 38]. For example, TartanVO [43] integrates deep learning layers to regress camera motion from optical flow, while DPVO [38] combines differentiable bundle adjustment with iterative correspondence updates. Dynamic environments introduce additional challenges, where moving objects can interfere with motion estimation. Improving VO robustness, some methods filter out dynamic regions with masks [2, 34], while others apply trajectory filtering to exclude moving elements [48, 7].
Dynamic Reconstruction.   Dynamic scene reconstruction aims to jointly recover geometry and motion in environments where objects and the camera are in motion. Traditional methods, such as SfM [32] and multi-view stereo (MVS) [33], assume static scenes, limiting their effectiveness in dynamic settings. Recent advances in monocular depth estimation leverage large-scale datasets and advanced neural architectures to infer depth from RGB sequences [4, 3, 29, 45, 18]. However, these methods often suffer from temporal inconsistencies across frames, leading to unstable depth predictions. To improve temporal coherence, video depth estimation techniques integrate strategies like depth consistency [25, 23] or diffusion priors [16]. Nevertheless, depth estimation alone often falls short of accurate 4D reconstruction. Joint optimization of the camera pose and depth is essential to avoid misalignment between depth predictions and camera motion. For example, RobustCVD [21] refines depth using 3D geometric constraints, while CasualSAM [47] finetunes a depth network with predicted motion masks. MonST3R [46] estimates geometry in dynamic scenes by leveraging the pointmap formulation [42].
@@ 3 BA-Track
The proposed BA-Track framework consists of three main components, see Fig. 3: (i) A learning-based front-end decouples camera-induced motion from observed motion via a 3D tracker. (ii) A bundle adjustment back-end recovers camera pose and sparse 3D geometry for both static and dynamic points using camera-induced motion. (iii) A lightweight refinement stage leverages the sparse tracks to produce dense and temporally consistent depth.
@@ 3.1 Stage I: Motion-Decoupled 3D Tracker
Classical correspondence-based methods struggle to recover the 3D structure of dynamic objects from monocular videos due to the epipolar constraint violation. As a result, previous work requires detecting and filtering out dynamic regions to ensure that only static points are used for camera tracking and reconstruction [2, 48, 7]. Here, we explore an alternative approach. Instead of discarding dynamic points, we infer the camera-induced component of their motion. As shown in Fig. 4, the observed 2D motion comprises camera-induced (static) and object-induced (dynamic) components. By accurately estimating the static component, dynamic points become pseudo-static in their local reference frame. Leveraging the static component, classical optimization, such as bundle adjustment, can operate seamlessly and require no special treatment of the dynamic points.
However, accurately estimating the camera-induced motion is non-trivial and requires strong motion priors. To effectively learn motion priors, we integrate off-the-shelf depth cues and develop a 3D tracking front-end. The tracker exploits the temporal context, while the additional depth cues enhance the tracker’s 3D reasoning and improve its capacity to estimate the camera-induced motion. Furthermore, we empirically observe that training a single network to estimate the camera-induced component is suboptimal. Instead, we employ two networks. The first predicts the observed, “total” motion, while the other isolates the dynamic component corresponding to the object-induced motion. We further discuss and empirically verify this design choice in Sec. 4.4.
@@
To extract relevant features for tracking, we pass each RGB image and its corresponding depth map through a CNN-based feature encoder [44], and get $C-$dimensional feature maps, $\mathbf{F}=\{F_{1},...,F_{S}\}$, where $F_{s}\in\mathbb{R}^{C\times H/4\times W/4}$. We then define a point feature vector $f(s)$ associated with a 2D point $\mathbf{x}$, which encapsulates appearance, spatial, and depth information from the RGB-D input at each timestep $s$.
@@
Motion Decoupling.   Our front-end consists of two transformer networks sharing the architecture with CoTracker [17]. The tracker $\mathcal{T}$ predicts the observed motion $X_{\text{total}}$, visualized in blue in Fig. 3. The dynamic tracker $\mathcal{T}_{\text{dyn}}$ is two times shallower than $\mathcal{T}$ for efficiency, and predicts the dynamic component $X_{\text{dyn}}$ of $X_{\text{total}}$. The tracker $\mathcal{T}$ also learns point visibility $v\in[0,1]^{S}$ and a static-dynamic label $m\in[0,1]$. Point visibility reflects occlusions, while the static-dynamic label indicates if the point belongs to a dynamic object.
@@
where $m$ acts as a gating factor. Observe that the motion of static points ($m=0$) coincides with the observed motion. Conversely, the camera-induced motion of dynamic points ($m=1$) results from subtracting their motion from the observed flow. By decoupling the motion into static and dynamic parts, we can effectively track dynamic points as if they were static in their local frame of reference. This novel mechanism allows the tracker to recover the 3D structure of dynamic objects with bundle adjustment in a seamless fashion, as we will elaborate in Sec. 3.2.
Training.   Under the hood, each transformer takes an initial 3D query location, extracts its point feature and iteratively updates the point location and the corresponding point feature. Each iteration leverages the local context of the query point, aggregated from the multi-scale feature correlations [17]. Additionally, we extend the context aggregation to the depth maps in order to embed the 3D context. The depth correlation is computed as the difference in depth values between the query point and its local neighbors. Let us denote the point location and its feature vector after $k$-th iteration as $(X^{(k)},\>f^{(k)})$ for tracker $\mathcal{T}$, and $(X_{\text{dyn}}^{(k)},f_{\text{dyn}}^{(k)})$ for the dynamic tracker $\mathcal{T}_{\text{dyn}}$. Correspondingly to Eq. 1, we obtain the static component for each iteration:
@@
We supervise $\mathcal{T}$ and $\mathcal{T}_{\text{dyn}}$ with ground-truth total 3D point trajectories and the static point trajectories from synthetic data. To improve the training convergence, we provide the supervision for every iteration of the tracker,
@@
To learn the visibility and the static-dynamic label, we use the binary cross-entropy, defined as
@@ 3.2 Stage II: Bundle Adjustment
We now employ our 3D tracker above for recovering the camera poses and a sparse 3D structure of dynamic scenes. Note that Sec. 3.1 exemplified tracking of a single query point, but our window-based 3D tracker can handle multiple query points within a temporal window. Here, we assume to have a set of 3D trajectories represented by their static 3D components $X_{\text{static}}$, as well as their per-frame visibility $v$ and the static-dynamic label $m$. To simplify the notation, we use $X$ to denote the static component $X_{\text{static}}$ in the following.
Given a video sequence of length $L>S$, we extract a set of $N$ query points $\mathbf{X}^{t}=(X^{t}_{1},\dots,X^{t}_{N})$ from each frame $t\in\{1,...,L\}$. Each query point $X^{t}_{n}=(u^{t}_{n},v^{t}_{n},d^{t}_{n})$ represents the 2D pixel coordinates $\mathbf{x}^{t}_{n}=(u^{t}_{n},v^{t}_{n})$ of the point and its initial depth $d^{t}_{n}=D_{t}[\mathbf{x}^{t}_{n}]$. Using our 3D tracker, we estimate a 3D trajectory of each query point $X^{t}_{n}$ within a local window of $S=2S^{\prime}+1$ frames. Specifically, we obtain the trajectory as $\mathbf{X}_{n}^{t}=[X_{n}^{t}(1),\dots,X_{n}^{t}(S)]$ defined over the window $I_{t-S^{\prime}:t+S^{\prime}}$. Using a sliding window approach, we collect the 3D trajectories of all query points over the full sequence. The result is a local 3D trajectory tensor $\mathbf{X}\in\mathbb{R}^{L\times N\times S\times 3}$.
Since our tracker can estimate the static component of each point trajectory, we can leverage the classical BA optimization framework [39] for camera pose estimation and depth refinement. This approach enhances the robustness of camera pose estimation, since it isolates the object motion and reduces the impact of dynamic points on pose accuracy. In our BA framework, the optimization variables include the per-frame camera poses $\{\mathbf{T}_{t}\in SE(3)\}$ and the refined depths of the query points $\mathbf{Y}\in\mathbb{R}^{L\times N\times 1}$. Each 3D trajectory $\mathbf{X}$ connects a query point $\mathbf{x}^{i}_{n}$ extracted from its source frame $i$ to its corresponding target locations in all other frames $j$.
@@
where $\left\|\cdot\right\|_{\rho}$ is the Huber loss and $\alpha$ balances the reprojection loss and depth consistency. The confidence weight $W^{i}_{j,n}$ encapsulates the visibility of point $\mathbf{x}_{n}^{i}$ in frame $j$ and its dynamic label by $W^{i}_{n}(j)=v^{i}_{n}(j)\cdot(1-m_{n}^{i}).$ We solve Eq. 8 efficiently using the Gauss-Newton method with Schur decomposition [39].
Category Method MPI Sintel [6] AirDOS Shibuya [30] Epic Fields [40]
@@
w/ intrinsics DROID-SLAM [37] 0.175 0.084 1.912 0.256 0.123 0.628 1.424 0.130 2.180
TartanVO [43] 0.238 0.093 1.305 0.135 0.030 0.249 1.490 0.121 1.548
DytanVO [34] 0.131 0.097 1.538 0.088 0.026 0.251 1.608 0.119 1.556
DPVO [38] 0.115 0.072 1.975 0.146 0.091 0.387 0.394 0.078 0.734
LEAP-VO [7] 0.089 0.066 1.250 0.031 0.111 0.124 0.486 0.071 1.018
w/o intrinsics Robust-CVD  [21] 0.360 0.154 3.443 - - - - - -
ParticleSfM  [48] 0.129 0.031 0.535 0.275 0.155 0.750 - - -
CasualSAM  [47] 0.141 0.035 0.615 0.209 0.202 0.620 - - -
MonST3R  [29] 0.108 0.042 0.732 (0.512) (0.075) (0.566) - - -
@@
Table 1: Camera pose evaluation results on Sintel [6], Shibuya [30], and Epic Fields [40] datasets. ()($\cdot$)( ⋅ ) denotes evaluation on sub-sequence due to memory constraints. Our method shows better results on ATE compared to other competitive baselines.

### v2

@@ Abstract
Traditional SLAM systems, which rely on bundle adjustment, struggle with the highly dynamic scenes commonly found in casual videos. Such videos entangle the motion of dynamic elements, undermining the assumption of static environments required by traditional systems. Existing techniques either filter out dynamic elements or model their motion independently. However, the former often results in incomplete reconstructions, while the latter can lead to inconsistent motion estimates. Taking a novel approach, this work leverages a 3D point tracker to separate camera-induced motion from the observed motion of dynamic objects. By considering only the camera-induced component, bundle adjustment can operate reliably on all scene elements. We further ensure depth consistency across video frames with lightweight post-processing based on scale maps. Our framework combines the core of traditional SLAM—bundle adjustment—with a robust learning-based 3D tracker. Integrating motion decomposition, bundle adjustment, and depth refinement, our unified framework, BA-Track, accurately tracks camera motion and produces temporally coherent and scale-consistent dense reconstructions, accommodating both static and dynamic elements. Our experiments on challenging datasets reveal significant improvements in camera pose estimation and 3D reconstruction accuracy. Project page: wrchen530.github.io/projects/batrack
@@ 1 Introduction
Figure 2: Dynamic scene reconstruction results on Shibuya [qiu2022airdos], DAVIS [khoreva2019video], and Aria Everyday Activities [lv2024aria] datasets. By leveraging sparse dynamic SLAM and global refinement, BA-Track achieves consistent dense 3D reconstructions across diverse dynamic scenes.
@@
With the increasing prevalence of casual videos, reconstructing scenes containing moving objects has become essential for applications like augmented reality and robotics. Traditional methods for structure-from-motion (SfM) [Agarwal:2011:BRD] and simultaneous localization and mapping (SLAM) [Engel:2014:LSD] excel at recovering camera poses and scene geometry. However, they leverage the epipolar constraint and thus apply only to static environments. Dynamic scenes with moving objects introduce ambiguities that violate the epipolar constraint. Resolving these ambiguities by analytical constraints or statistical outlier filtering alone is infeasible due to the complexity and diversity of real-world object motion. This challenge motivates a hybrid approach that combines traditional optimization with deep priors. Existing hybrid methods often address dynamic elements by detecting and removing them from optimization, leading to incomplete reconstructions or missing details associated with moving objects [bescos2018dynaslam, zhao2022particlesfm]. Another approach is to rely on geometric cues, such as monocular depth models for per-frame reconstruction, but this requires additional optimization to achieve a globally aligned 3D reconstruction [kopf2021robust]. This proves challenging in practice due to the inconsistency between estimated camera pose and depth priors across different frames, which can result in misaligned depth maps when integrating them into a coherent 3D scene.
@@
We propose BA-Track, a framework that jointly reconstructs static and dynamic scene geometry from casual video sequences (see Fig. 1). Following traditional SfM and SLAM approaches, BA-Track extends the time-honored bundle adjustment (BA) [Triggs:1999:BAA] to dynamic scene reconstruction. The novel insight is to integrate a learning-based 3D tracker to decouple the motion of dynamic elements from the observed motion. By disentangling the camera-induced motion component of dynamic objects, BA-Track reconstructs 3D trajectories of dynamic points as if they were stationary relative to their local reference frame. As a result, the epipolar constraint becomes applicable to all points, dynamic or static, enabling robust operation of bundle adjustment.
@@
To effectively decouple motion, we incorporate monocular depth priors with a 3D tracker. The tracker learns valuable 3D priors that help distinguish between camera- and object-based motion. Although such learning-based priors prove extremely helpful for reconstructing dynamic scenes, the scale of the depth maps exhibits temporal and spatial inconsistency. To address this issue, we develop a global refinement module, which leverages the accurate but sparse geometry from BA to refine the dense depth maps. As illustrated in Fig. 2, BA-Track achieves compelling reconstruction quality of both static and dynamic scene elements.
@@
In summary, our framework consists of three components: (1) a robust 3D tracking front-end, which decouples camera-induced motion from observed (total) motion; (2) a bundle adjustment back-end for accurate pose and depth estimation, which leverages camera-induced motion for all points; and (3) global refinement, to achieve dense, scale-consistent depth. Our learnable pipeline extends epipolar geometry to dynamic scene reconstruction, leveraging the power of established optimization techniques. Our experiments demonstrate significant improvements in camera motion estimation and 3D scene reconstruction, even in challenging scenes.
@@ 2 Related Work
Point Tracking.  Point tracking techniques estimate the 2D positions of one or more query points across a sequence of images. Revisiting the idea of video representation as moving particles [sand2008particle], PIPs [harley2022particle] takes a learning-based approach with a feed-forward network that iteratively refines point trajectories using multi-scale local features. TAP-Net [doersch2022tap] takes a different direction by leveraging global correlations to directly compute trajectories, thereby eliminating the need for iterative refinement. Subsequent research has explored various techniques to improve accuracy and efficiency, including inter-track relationships [karaev2023cotracker], temporal refinements [doersch2023tapir], nearest-neighbor interpolation [le2024dense], and bidirectional 4D correlation [cho2025local]. Moving beyond 2D tracking, recent methods such as SpatialTracker [xiao2024spatialtracker] and SceneTracker [bescos2018dynaslam] elevate the task into 3D space by incorporating depth priors.
Our approach extends 3D tracking by introducing a novel motion decoupling strategy that separates camera motion from object motion, enabling more effective integration with bundle adjustment.
@@
Monocular Visual Odometry.  Monocular visual odometry (VO) estimates camera motion using RGB image sequences. Traditional methods can be divided into feature-based and direct approaches. Feature-based methods detect and match keypoints across frames, using epipolar geometry and bundle adjustment for pose estimation [davison2007monoslam, mur2015orb]. Direct methods bypass keypoints, instead optimizing photometric consistency across frames using full-image information [Engel:2014:LSD, gao2018ldso, engel2017direct]. Recent advances have shifted towards learning-based approaches that employ neural networks to enhance feature extraction and pose estimation [bloesch2018codeslam, teed2018deepv2d, teed2021droid, tang2018ba, wang2021tartanvo, teed2024deep, wimbauer2025anycam]. For example, TartanVO [wang2021tartanvo] regresses motion from optical flow using deep networks, while DPVO [teed2024deep] integrates differentiable bundle adjustment with iterative correspondence updates. Dynamic environments introduce additional challenges, as moving objects can interfere with motion estimation. To improve robustness, some methods filter dynamic regions using masks [bescos2018dynaslam, shen2023dytanvo], while others apply trajectory filtering to discard motion outliers [zhao2022particlesfm, Chen:2024:LEAP].
Dynamic Scene Reconstruction.  Dynamic scene reconstruction aims to jointly recover geometry and motion in environments with moving objects. Traditional methods such as SfM [schonberger2016structure] and MVS [schonberger2016pixelwise] assume static scenes, limiting their applicability in dynamic settings. Regression-based approaches leverage large-scale training to estimate depth from single RGB images [birkl2023midas, bhat2023zoedepth, piccinelli2024unidepth, yang2024depthanyv2, ke2024repurposing], while video-based extensions further improve temporal coherence by incorporating multi-frame cues, depth consistency, or diffusion priors [luo2020consistent, li2021enforcing, feng2022disentangling, hu2024depthcrafter]. More recent approaches extend beyond depth to regress scene coordinates, modeling dynamic scenes through per-frame point maps [zhang2024monst3r, wang2025continuous, sucar2025dynamic, feng2025st4rtrack]. To enforce geometric consistency, several methods combine learning with optimization, jointly refining scene structure and camera motion [kopf2021robust, zhang2022structure, zhang2024monst3r, li2025megasam]. For example, MonST3R [zhang2024monst3r] uses a point map representation and refines global poses by filtering dynamic regions with an optical flow-based motion mask. TracksTo4D [kasten2024fast], closely related to our work, reconstructs dynamic scenes from 2D point trajectories by directly regressing point cloud bases and camera poses. In contrast, our method decouples static and dynamic point motions and jointly optimizes them via bundle adjustment, achieving better robustness.
@@ 3 BA-Track
BA-Track consists of three stages (see Fig. 3): (i) A learning-based front-end decouples camera-induced motion from observed motion via a 3D tracker. (ii) A bundle adjustment back-end recovers camera pose and sparse 3D geometry for both static and dynamic points using camera-induced motion. (iii) A lightweight refinement stage leverages the sparse tracks to produce dense and temporally consistent depth.
@@ 3.1 Stage I: Motion-Decoupled 3D Tracker
Classical correspondence-based methods struggle to recover the 3D structure of dynamic objects from monocular videos due to violations of the epipolar constraint. As a result, previous work requires detecting and filtering out dynamic regions to ensure that only static points are used for camera tracking and reconstruction [bescos2018dynaslam, zhao2022particlesfm, Chen:2024:LEAP]. Here, we explore an alternative approach. Instead of discarding dynamic points, we infer the camera-induced component of their motion. As shown in Fig. 4, the observed 2D motion comprises camera-induced (static) and object-induced (dynamic) components. By accurately estimating the static component, dynamic points become pseudo-static in their local reference frame. Leveraging the static component, classical optimization, such as bundle adjustment, can operate seamlessly and requires no special treatment of dynamic points.
However, accurately estimating the camera-induced motion is non-trivial and requires strong motion priors. To effectively learn motion priors, we integrate off-the-shelf depth cues and develop a 3D tracking front-end. The tracker exploits the temporal context, while the additional depth cues enhance the tracker’s 3D reasoning and improve its capacity to estimate the camera-induced motion. Furthermore, we empirically observe that training a single network to estimate the camera-induced component is suboptimal. Instead, we employ two networks: the first predicts the observed, “total” motion, while the other predicts the dynamic component corresponding to the object-induced motion. We further discuss and empirically verify this design choice in Sec. 4.4.
@@
To extract relevant features for tracking, we pass each RGB image and its corresponding depth map through a CNN-based feature encoder [xiao2024spatialtracker], and obtain $C$-dimensional feature maps, $\mathbf{F}=\{F_{1},\dots,F_{S}\}$, where $F_{s}\in\mathbb{R}^{C\times H/4\times W/4}$. We then define a point feature vector $f(s)$ associated with a 2D point $\mathbf{x}$, which encapsulates appearance, spatial, and depth information from the RGB-D input at each timestep $s$.
@@
Motion Decoupling.  Our front-end consists of two transformer networks sharing the architecture with CoTracker [karaev2023cotracker]. The tracker $\mathcal{T}$ predicts the observed motion $X_{\text{total}}$, visualized in blue in Fig. 3. The dynamic tracker $\mathcal{T}_{\text{dyn}}$ is two times shallower than $\mathcal{T}$ for efficiency, and predicts the dynamic component $X_{\text{dyn}}$ of $X_{\text{total}}$. The tracker $\mathcal{T}$ also learns point visibility $v\in[0,1]^{S}$ and a static-dynamic label $m\in[0,1]$. Point visibility reflects occlusions, while the static-dynamic label indicates if the point belongs to a dynamic object.
@@
where $m$ acts as a gating factor. Observe that the motion of static points ($m=0$) coincides with the observed motion. Conversely, the camera-induced motion of dynamic points ($m=1$) results from subtracting their motion from the observed flow. By decoupling the motion into static and dynamic parts, we can effectively track dynamic points as if they were static in their local frame of reference. This novel mechanism allows the tracker to recover the 3D structure of dynamic objects with bundle adjustment in a seamless fashion, as we elaborate in Sec. 3.2.
Training.  Under the hood, each transformer takes an initial 3D query location, extracts its point feature, and iteratively updates the point location and the corresponding point feature. Each iteration leverages the local context of the query point, aggregated from the multi-scale feature correlations [karaev2023cotracker]. We further incorporate depth map features into the context aggregation to embed 3D context. Let us denote the point location and its feature vector after the $k$-th iteration as $(X^{(k)},\>f^{(k)})$ for tracker $\mathcal{T}$, and $(X_{\text{dyn}}^{(k)},f_{\text{dyn}}^{(k)})$ for the dynamic tracker $\mathcal{T}_{\text{dyn}}$. Correspondingly to Eq. 1, we obtain the static component for each iteration:
@@
We supervise $\mathcal{T}$ and $\mathcal{T}_{\text{dyn}}$ with ground-truth total 3D point trajectories and the static point trajectories from synthetic data. To improve training convergence, we provide supervision for every iteration of the tracker,
@@
To learn the visibility and the static-dynamic label, we use binary cross-entropy, defined as
@@ 3.2 Stage II: Bundle Adjustment
We now employ our 3D tracker above to recover the camera poses and a sparse 3D structure of dynamic scenes. Note that Sec. 3.1 exemplified tracking of a single query point, but our window-based 3D tracker can handle multiple query points within a temporal window. Here, we assume to have a set of 3D trajectories represented by their static 3D components $X_{\text{static}}$, as well as their per-frame visibility $v$ and the static-dynamic label $m$. To simplify the notation, we use $X$ to denote the static component $X_{\text{static}}$ in the following.
Given a video sequence of length $L>S$, we extract a set of $N$ query points $\mathbf{X}^{t}=(X^{t}_{1},\dots,X^{t}_{N})$ from each frame $t\in\{1,\dots,L\}$. Each query point $X^{t}_{n}=(u^{t}_{n},v^{t}_{n},d^{t}_{n})$ represents the 2D pixel coordinates $\mathbf{x}^{t}_{n}=(u^{t}_{n},v^{t}_{n})$ of the point and its initial depth $d^{t}_{n}=D_{t}[\mathbf{x}^{t}_{n}]$. Using our 3D tracker, we estimate a 3D trajectory of each query point $X^{t}_{n}$ within a local window of $S=2S^{\prime}+1$ frames. Specifically, we obtain the trajectory as $\mathbf{X}_{n}^{t}=[X_{n}^{t}(1),\dots,X_{n}^{t}(S)]$ defined over the window $I_{t-S^{\prime}:t+S^{\prime}}$. Using a sliding window approach, we collect the 3D trajectories of all query points over the full sequence. The result is a local 3D trajectory tensor $\mathbf{X}\in\mathbb{R}^{L\times N\times S\times 3}$.
Since our tracker can estimate the static component of each point trajectory, we can leverage the classical BA optimization framework [Triggs:1999:BAA] for camera pose estimation and depth refinement. This approach enhances the robustness of camera pose estimation, since it isolates the object motion and reduces the impact of dynamic points on pose accuracy. In our BA framework, the optimization variables include the per-frame camera poses $\{\mathbf{T}_{t}\in SE(3)\}$ and the refined depths of the query points $\mathbf{Y}\in\mathbb{R}^{L\times N\times 1}$. Each 3D trajectory $\mathbf{X}$ connects a query point $\mathbf{x}^{i}_{n}$ extracted from its source frame $i$ to its corresponding target locations in all other frames $j$.
@@
where $\left\|\cdot\right\|_{\rho}$ is the Huber loss and $\alpha$ balances the reprojection loss and depth consistency. The confidence weight $W^{i}_{j,n}$ encapsulates the visibility of point $\mathbf{x}_{n}^{i}$ in frame $j$ and its dynamic label by $W^{i}_{n}(j)=v^{i}_{n}(j)\cdot(1-m_{n}^{i})$. We solve Eq. 8 efficiently using the Gauss-Newton method with Schur decomposition [Triggs:1999:BAA].
Category Method MPI Sintel [Butler:2012:Sintel] AirDOS Shibuya [qiu2022airdos] Epic Fields [tschernezki2024epic]
@@
w/ intrinsics DROID-SLAM [teed2021droid] 0.175 0.084 1.912 0.256 0.123 0.628 1.424 0.130 2.180
TartanVO [wang2021tartanvo] 0.238 0.093 1.305 0.135 0.030 0.249 1.490 0.121 1.548
DytanVO [shen2023dytanvo] 0.131 0.097 1.538 0.088 0.026 0.251 1.608 0.119 1.556
DPVO [teed2024deep] 0.115 0.072 1.975 0.146 0.091 0.387 0.394 0.078 0.734
LEAP-VO [Chen:2024:LEAP] 0.089 0.066 1.250 0.031 0.111 0.124 0.486 0.071 1.018
w/o intrinsics Robust-CVD [kopf2021robust] 0.360 0.154 3.443
ParticleSfM [zhao2022particlesfm] 0.129 0.031 0.535 0.275 0.155 0.750
CasualSAM [zhang2022structure] 0.141 0.035 0.615 0.209 0.202 0.620
MonST3R [piccinelli2024unidepth] 0.108 0.042 0.732 (0.512) (0.075) (0.566)
@@
Table 1: Camera pose evaluation results on Sintel [Butler:2012:Sintel], Shibuya [qiu2022airdos], and Epic Fields [tschernezki2024epic] datasets. $(\cdot)$ denotes evaluation on sub-sequence due to memory constraints. Our method shows better results on ATE compared to other competitive baselines.